# Is Limited Participant Diversity Impeding EEG-based Machine Learning?

### v1

@@ 2.3 Model Comparison
2.4 Participant Distribution Shifts
So far, we only varied the number of participants ($n$) and kept the number of segments per participant ($m$) fixed. Improvements in performance were therefore expected irrespective of the presence of participant distribution shifts. Figure 3 now illustrates the performance of the TCN model when controlling for both $m$ and $n$. Note that the diagonals extending from the top left to the bottom right of each heatmap correspond to fixed overall sample sizes. Moving from the bottom right along the diagonal towards the top left corresponds to the same overall sample size where data was sampled from a higher number of individual participants, i.e. higher participant diversity.
Focusing first on the normality prediction task on the TUAB dataset, we observe that increasing the amount of data for a fixed participant count (moving along the horizontal axis) did not yield any substantial performance improvements. However, performance was highly dependent on the participant count (moving along the vertical axis). Even for a fixed sample size, higher participant diversity yielded significant improvements (moving diagonally towards the top left). The results for the dementia prediction task on the CAUEEG dataset exhibit a similar pattern. Linking back to the theoretical analysis in Section 4.1, this suggests a substantial distribution shift between participants such that the participant count $n$ dominated model generalisation.
For the sleep staging task on the PhysioNet dataset the pattern is different. While higher participant diversity still yielded some improvements in performance (moving diagonally towards the top left), substantial improvements were also achieved by increasing the sample size for a fixed participant count (moving along the horizontal axis). In fact, with the maximum number of samples per participant, training data from a small cohort of just 25 participants was sufficient to match the performance achieved with training data from 400 participants (with 5 segments per participant).
Results for the mAtt model revealed the same patterns. For LaBraM, performance improvements for additional samples at a fixed participant count did not diminish as drastically, which is likely due to its lower sample efficiency as hypothesised in Section 2.3. Heatmaps for both models can be found in Section A.1.
2.5 Effectiveness of EEG Data Augmentation
Figure 4 illustrates the effect of different data augmentations chosen based on their promising performance in prior work (AmplitudeScaling – channel-wise multiplication with a random factor between 0.5 and 2, FrequencyShift – random shift in the frequency domain of up to 0.3 Hz, PhaseRandomization – randomisation of the phase of each Fourier coefficient; see Section 4.4).
Figure 4: Impact of data augmentations. Each boxplot shows pairwise accuracy differences between augmented and unaugmented training (positive means augmentation improved performance) across all combinations of participant counts ($n$) and segments per participant ($m$). Results for only a single seed are shown for better visibility. AS = AmplitudeScaling, FS = FrequencyShift, PR = PhaseRandomization. On PhysioNet, the LaBraM model benefitted from FS and PR, whereas PR decreased performance for TCN and mAtt. Other than that, augmentations did not lead to consistent improvements in performance.
With one exception, the impact of the different augmentations was generally small and inconsistent. What stood out, however, was a systematically detrimental effect of the PhaseRandomization augmentation on PhysioNet for the TCN and mAtt models, whereas performance for the LaBraM model was improved. LaBraM also benefitted from the FrequencyShift augmentation on PhysioNet. In terms of absolute performance, LaBraM performed substantially worse on PhysioNet without data augmentation (see Figure 2), such that with data augmentation, all models performed about equally well. The observed performance improvements for LaBraM can therefore likely be attributed to its greater reliance on large amounts of data, which is partially offset by augmentation. Decreased performance in the smaller and more data-efficient models suggests that PhaseRandomization destroys useful information.
The PhaseRandomization augmentation randomises the phase of each Fourier coefficient and is applied channel-wise. It therefore preserves the power distribution across different frequencies, but completely destroys wave morphologies and cross-channel correlations. Given that certain wave forms are highly informative of the sleep stage (e.g. K-complexes are known to occur during NREM 2 sleep), the performance decreases are certainly plausible.
2.6 Effectiveness of Self-supervised Pre-training
Figure 5 shows a comparison of the LaBraM model performance with and without pre-training. LaBraM was designed as an EEG foundation model and pre-trained in an unsupervised manner (i.e. without using labels) on a diverse collection of 16 datasets ranging from brain computer interface (BCI) datasets to emotion recognition and resting state data [jiang_large_2023]. In total, these datasets contained around 500 participants. After pre-training, the model was fine-tuned with the labelled training data of the corresponding task. The baseline model (without pre-training) was directly trained on the labelled training data from scratch. See Section 4.5 for details.
Pre-training consistently improved model performance across all data regimes, except when the amount of downstream fine-tuning data was most severely limited. Comparing columns in the left two heatmaps showing accuracies for the baseline (trained from scratch) and pre-trained model respectively, the effect of pre-training looks like a boost in the amount of data per participant: with only 10-20 segments per participant, the pre-trained model achieved the same performance as the baseline model with the maximum amount of 320 segments per participant. However, pre-training also improved performance when the participant count was low and it continued to provide value even when the maximum amount of data was used. Consistent results were observed for dementia prediction on the CAUEEG dataset and sleep staging on PhysioNet (see Section A.3).
In Section 2.2, we observed that the LaBraM model required substantially more data to achieve the same performance as the mAtt and TCN models. With pre-training, however, the much larger and more data-demanding LaBraM model was able to match or even outperform them. This can be seen by comparing the heatmaps in Figure 5 to the corresponding heatmap on TUAB of the TCN model in Figure 3. Alternatively, Section A.2 shows a version of Figure 2 with additional lines for the pre-trained LaBraM model.
3 Discussion
The present study demonstrates that the generalisability of EEG-based machine learning models to data from new participants is highly task-dependent. We observe that good model performance can be achieved with a relatively small number of participants in sleep staging, whereas performance in EEG normality prediction and dementia diagnosis was bottlenecked by the participant count. While our findings, therefore, do not admit a general conclusion—participant diversity is neither always the limiting factor nor universally unimportant—they highlight the importance of understanding its role in specific contexts. We recommend the conduction of pilot studies or review of prior application-specific work to inform the extent to which participant diversity should be prioritised in data collection and methodological machine learning work.
[…]
Self-supervised pre-training consistently improved model performance across all data regimes. These results are particularly promising considering the broad potential for further improvements: This study benchmarked LaBraM Base, the smallest version of the model for which weights were made publicly available by the authors. Further improvements are thus likely achievable with increased model size as indicated by comparisons of the model versions in the original publication [jiang_large_2023] as well as generally observed trends in self-supervised learning [kaplan_scaling_2020]. Moreover, the data used to pre-train LaBraM is still relatively limited with a total participant count of around 500—less than a third of the maximum number of participants used in our evaluations on the TUAB dataset. The existing body of literature on self-supervised learning strongly suggests that more data would result in better performance [kaplan_scaling_2020]. Since pre-training does not require labelled data, the creation of a larger pre-training datasets is comparatively easy. Prior work has also shown that self-supervised learning can be used to enforce approximate invariances [ericsson2022why], which can help circumvent the poor dependence of overfitting behaviour on the number of participants in the training data [gouk_limitations_2024].
Beyond augmentations and self-supervised learning, refined preprocessing methods could provide another approach to dealing with limited EEG data by improving data quality instead of quantity. There is a rich literature on EEG preprocessing [debnath_maryland_2020, jas_autoreject_2017, pion-tonachini_iclabel_2019], although most methods were designed for classical EEG analysis and it is unclear how well they translate to ML-based EEG applications. Some work suggests that even with large datasets, machine learning models still benefit in particular from the removal of high-amplitude artefacts [gjolbye_speed_2024, bomatter_machine_2024].
Finally, future research should attempt to determine the nature of participant distribution shifts. We observed a large participant count to be critical for model performance in EEG normality prediction and EEG-based dementia diagnosis, and we can see several plausible underlying reasons. On the one hand, anatomical individual differences could impede the generalisability of the observed patterns of scalp activity measured with EEG. On the other hand, the distribution shift might also be explained by variability in how the target presents in different participants. For instance, in the EEG normality prediction task, there are numerous reasons why an EEG can be abnormal, ranging from focal epileptiform discharges to diffuse slowing. A model that is trained on data from only a small cohort of participants will therefore not only be exposed to little demographic and anatomical variation but also to a limited number of target (or disease) phenotypes. This could be particularly relevant, e.g., in the context of neurodegenerative or psychiatric diseases, where substantial individual differences in symptomatology and neural correlates have been described [young_uncovering_2018, buch_dissecting_2021].
4 Methods
4.1 Experimental Design and Model Training
Each dataset was split into train, validation, and test splits such that participants did not overlap between the splits. To study the impact of the amount and participant diversity of the training data, each experiment consisted of a grid of trials where we varied the number of participants and the number of segments per participant. Training data for each configuration was chosen through random sub-sampling from the train split (see Figure 1). To obtain uncertainty estimates, each trial was repeated multiple times with different random seeds for the data sub-sampling.
Models were trained using AdamW (betas=(0.9, 0.999),
weight_decay=0.01), global gradient norm clipping (max_norm = 1.0), and a cross-entropy loss. The learning rate was set to 1e-3 for mAtt and TCN and tuned for LaBraM (see Section 4.5 and Table 1 for details and the used learning rates respectively). For the CAUEEG dataset, we used class weights in the loss function to account for class imbalance. Since the amount and diversity of the training data differed substantially between trials, early stopping (on the fixed validation set) was used to prevent overfitting, while allowing sufficient training time if the amount of data was larger. Specifically, the validation loss was evaluated every 500 batches and a patience of 5 was used to trigger early stopping. If early stopping was triggered during training, the best model was restored for evaluation.
Finally, all trained models were evaluated on the same test data. We ran each experiment for different models (Section 4.3), datasets, and tasks (Section 4.2). To study different strategies to deal with limited data, we repeated the same procedure for each studied method (augmentation: see Section 4.4, self-supervised learning: see Section 4.3.3). The research code for our experiments is publicly available online [^2] https://github.com/bomatter/participant-diversity-paper.
4.2 Data and Tasks
4.2.1 TUAB
The Temple University Hospital Abnormal (TUAB) EEG Corpus [lopez_automated_2015, diego_automated_2017] is a demographically balanced and curated subset of the TUH EEG Corpus [obeid_temple_2016]. The data was collected in a hospital environment and interpreted by board certified neurologists. The resulting medical reports contain a decision of whether the EEG was found to be normal or abnormal. The assignment of this label is standardised with clear criteria about what constitutes a normal or abnormal EEG [rubin_clinical_2021]. However, it is worth noting that an EEG can be abnormal for many different underlying reasons. Labels for the TUAB dataset were extracted automatically from the medical reports using a natural language processing approach and subsequently verified manually by a team of students. The full dataset contains more than 2300 participants with most recording durations in the range of 15-25 minutes. For the current study, we used version 3.0.1 of the dataset.
4.2.2 CAUEEG
The Chung-Ang University Hospital EEG (CAUEEG) dataset [kim_deep_2023] comprises data from 1379 participants with most recording durations in the range of 5 - 20 minutes. In the current study, we used a subset of data from 1122 participants for which diagnostic labels for the categories normal, mild cognitive impairment (MCI), or dementia were available. Diagnostic labels were assigned by neurologists based on a neuropsychological examination.
4.2.3 PhysioNet
The dataset referred to as PhysioNet in this study corresponds to the data used in the PhysioNet/Computing in Cardiology Challenge 2018 [ghassemi_you_2019, goldberger_physiobank_2000]. While the official challenge was focused on the classification of arousal regions, we focused on sleep staging here. The dataset comprises polysomnographic (PSG) recordings along with sleep stage annotations for non-overlapping 30-second epochs following AASM standards. While the PSG recordings contained additional measurements such as electrooculography (EOG), electromyography (EMG), or airflow and oxygen saturation, only the six EEG channels were used in this study. Furthermore, we only used the training data of version 1.0.0 of the dataset, as annotations for the official test split were not made publicly available.
4.2.4 Preprocessing
Preprocessing of the data comprised the following steps: Band-pass filtering to 1-50 Hz and resampling to a sampling frequency of 100 Hz, cropping (to 15 minutes for TUAB, omitting the first 30 seconds due to frequent strong artefacts at the beginning of recordings; to 5 minutes for CAUEEG), selection of a fixed channel subset (the 19 10-20 channels ”Fp1”, ”Fp2”, ”F7”, ”F3”, ”Fz”, ”F4”, ”F8”, ”T7”, ”C3”, ”Cz”, ”C4”, ”T8”, ”P7”, ”P3”, ”Pz”, ”P4”, ”P8”, ”O1”, ”O2” for TUAB and CAUEEG; the 6 channels ”F3”, ”F4”, ”C3”, ”C4”, ”O1”, ”O2” for PhysioNet), re-referencing to average reference, and segmentation into non-overlapping epochs (2 seconds for TUAB and CAUEEG; 30 seconds for PhysioNet).
For compatibility with the pre-trained LaBraM model, the following adjustments were made for the training and fine-tuning of LaBraM: amplitudes scaled to units of 0.1 mV, sampling frequency of 200 Hz, filtering to 0.1-75 Hz with a notch filter at 50 Hz (see [jiang_large_2023]). We note that the notch filter was not matched to the line noise of the data in the original publication (e.g. LaBraM was also evaluated on TUAB, where line noise is at 60 Hz). For compatibility with the pre-trained weights, we did not correct this potentially suboptimal setting. Furthermore, since LaBraM only supports segment lengths up to 16 seconds without modification (due to the implementation of the temporal positional encodings), we used 15 second epochs instead of 30 second epochs for PhysioNet.
4.3 Models
4.3.1 TCN
[…]
4.3.2 mAtt
The manifold attention network (mAtt) is a geometric deep learning model developed for EEG data [pan_matt_2022]. Briefly, the architecture starts with two convolutional layers and then encodes segments in the input EEG into a sequence of covariance matrices. The diagonals in these covariance matrices correspond to the signal power in each channel. The architecture then makes use of the fact that the covariance matrices are symmetric positive definite (SPD) through a manifold attention module, where instead of dot product attention, an approximation of the Riemannian distance is used. The output is then transformed back to Euclidian space and passed through a fully connected layer to get classifications. We adapted the code from the official mAtt repository (https://github.com/CECNL/MAtt) and used 100 kernels in conv1, 50 kernels in conv2 with a kernel size of 11, and an embedding size of 25 for the SPD transforms.
4.3.3 LaBraM
[…]
4.4 Augmentations
4.4.1 Amplitude Scaling
[…]
Intuitively, amplitude scaling should make a model more robust to small changes in amplitude. Moreover, it could simulate some variability in electrode placement, skin conductance, or even individual differences between participants (e.g. skull thickness). On the other hand, we note that while AmplitudeScaling was intended to preserve the interpretation of the signal, overall amplitude and differences in amplitude between EEG channels (e.g. asymmetries between corresponding channels on the left and right hemisphere) can provide clinically relevant information [rubin_clinical_2021].
4.4.2 Frequency Shift
[…]
The motivation for FrequencyShift are individual differences in peak frequencies. Such differences are well established in EEG research and can be associated with demographic variables. For example, alpha peak frequency is known to change with age [aurlien_eeg_2004, chiang_age_2011, cellier_development_2021] and female participants were observed to have higher peak frequencies than male participants [aurlien_eeg_2004, chiang_age_2011]. At the same time, frequency shifts can be clinically relevant (e.g. slowing of alpha, theta, or delta activity may reflect abnormalities [rubin_clinical_2021]). Depending on the task, this augmentation could therefore also have undesirable effects and the maximum frequency shift is likely an important parameter.
4.4.3 Phase Randomisation
[…]
Intuitively, PhaseRandomisation keeps the signal’s power spectrum intact while wave form information is completely lost. This may help the model to leverage frequency information and it enforces invariance to wave morphologies. Since the power spectrum varies considerably across participants, this augmentation is likely to immitate more data from the same participants, and it might be less effective in increasing participant diversity. Furthermore, there are many examples of informative wave morphologies ranging from sleep spindles and K-complexes in sleep staging [thomas_atlas_2023] to sharp waves and spikes in the diagnosis of epilepsy [rubin_clinical_2021]. The enforced invariance to wave morphology could therefore be undesirable.
4.5 Pre-training
The LaBraM model (see Section 4.3.3) was developed as a foundation model for EEG data and pre-trained to predict discrete tokens, previously learned by a separate tokeniser that was itself trained to predict the Fourier spectrum (see the original publication for details [jiang_large_2023]). The model was pre-trained in an unsupervised manner on a collection of 16 datasets with a total of around 500 participants, and the weights for the smallest version (LaBraM Base) were made publicly available by the authors. We used these weights for experiments with the pre-trained LaBraM model.
The model was then fine-tuned using the labelled training data of the corresponding task. We fine-tuned the entire architecture without freezing any layers and determined the learning rate through hyperparameter optimisation. Specifically, we evaluated a grid of learning rates [1e-7, 5e-7, 1e-6, …, 1e-2] for each dataset (TUAB, CAUEEG, PhysioNet) using the respective training and validation splits. To ensure that performance differences observed between the pre-trained and baseline (trained from scratch) versions can be attributed to the pre-training rather than tuning of the learning rate, we also tuned the learning rate for the baseline in the same way. Table 1 shows the learning rates used for the different conditions.
Dataset From Scratch Fine-tuning
TUAB $41e-$ 4 $61e-$ 6
CAUEEG $31e-$ 3 $65e-$ 6
PhysioNet $31e-$ 3 $51e-$ 5
Table 1: Learning rates for the LaBraM model.
4.6 Statistical Analysis
As outlined in Section 4.1, each trial (i.e. each combination of a given number of participants and number of segments per participant) was repeated multiple times with different random seeds for the subsampling of the training data. This allowed us to report average accuracies along with the standard error of the mean across seeds. The baseline experiments were repeated with 25 seeds and the pre-training experiments were repeated with 25 seeds on TUAB and CAUEEG and with 5 seeds on PhysioNet. For the augmentations, we report pairwise comparisons for a single seed. On the PhysioNet dataset, where considerably longer recordings were available for some participants, we assessed conditions with up to 640 segments per participant. However, since this number of segments is not available for all participants in the dataset, experiments failed for some of the seeds, which is also the reason for the missing tiles in the top right corner of the PhysioNet results heatmaps. Rather than only reporting the results up to 40 segments per participant, up to which point we have complete results for all seeds, we included trials with failed seeds as they provide interesting and unbiased additional information.
[…]
5 Acknowledgements
Appendix A Appendix
A.1 Accuracy Heatmaps for mAtt and LaBraM
@@
A.2 Ablation: LaBraM-compatible Preprocessing
Data for the training and fine-tuning of the LaBraM model was preprocessed separately to ensure compatibility with the official pre-trained weights. On the PhysioNet dataset, this also meant that the segment length had to be reduced from 30s to 15s. Since LaBraM’s performance on PhysioNet was particularly low compared to the other models, we performed an ablation study to see how much of the performance gap should be attributed to the different preprocessing rather than to architectural differences. Specifically, we trained the TCN model using the same preprocessing that was used for the LaBraM model. Figure 8 shows that performance decreased (light blue line vs dark blue line) but was still higher than the LaBraM baseline performance (light green line). With pre-training, LaBraM achieved comparable performance to the TCN model when the same preprocessing was used (dark blue and dark green lines). Overall, this suggests that the different preprocessing explains part of the performance gap, but not all of it. When models were trained from scratch, TCN still achieved better performance.
A.3 Pre-training Results on CAUEEG and PhysioNet
@@
A.4 Scaling Behaviour Plots with Pre-trained LaBraM Model
Figure 11: Scaling behaviour with pre-trained LaBraM model. Lines show the average accuracy across seeds with the standard error of the mean. Pre-training consistently improved LaBraM’s performance and allowed it to achieve comparable or even superior results to the other models. The lower performance on the PhysioNet dataset is due to differences in preprocessing (see Section A.2).

### v2

@@ A.5 Model Comparison
A.6 Ablation: LaBraM-compatible Preprocessing
Data for the training and fine-tuning of the LaBraM model was preprocessed separately to ensure compatibility with the official pre-trained weights. On the PhysioNet dataset, this also meant that the segment length had to be reduced from 30s to 15s. Since LaBraM’s performance on PhysioNet was particularly low compared to the other models, we performed an ablation study to see how much of the performance gap should be attributed to the different preprocessing rather than to architectural differences. Specifically, we trained the TCN model using the same preprocessing that was used for the LaBraM model. Figure˜6 shows that performance decreased (light blue line vs dark blue line) but was still higher than the LaBraM baseline performance (light green line). With pre-training, LaBraM achieved comparable performance to the TCN model when the same preprocessing was used (dark blue and dark green lines). Overall, this suggests that the different preprocessing explains part of the performance gap, but not all of it. When models were trained from scratch, TCN still achieved better performance.
A.7 Control Experiment: Contiguous Segment Subsampling
In our main experiments, segments were subsampled uniformly at random. However, in practice, collecting fewer segments would usually correspond to shorter recordings, which are less likely to capture a wide range of long-term non-stationarities such as drifts, participant fatigue, or changes in mental state. Uniform subsampling may therefore artificially increase segment diversity compared to realistic data collection.
To examine whether this affects our conclusions, we performed a control experiment using contiguous subsampling, where training segments are drawn from consecutive blocks rather than uniformly. Specifically, for an epoched EEG with $N$ non-overlapping fixed-length segments and a desired number of segments $m$, we uniformly sample a random starting index $s\in[0,N-m]$. We then extract a contiguous block of $m$ segments: $\text{segments}[s:s+m]$.
Results are shown in Figure˜7, which compares performance with uniform versus contiguous subsampling for normality prediction on TUAB with the TCN model. Patterns are consistent across both conditions, indicating that the conclusions are not affected by the choice of the subsampling strategy.
Figure 7: Uniform versus contiguous segment subsampling. Normality prediction accuracy on TUAB with the TCN model, averaged over 5 random seeds along with the standard error of the mean.
A.8 Accuracy Heatmaps for mAtt and LaBraM
@@
A.9 No Association between Sample Size and Accuracy Improvements with Augmentations
Table 2: Spearman correlations between sample size and accuracy difference to baseline performance with different augmentations. No associations between accuracy improvements and sample size were observed. $p$-values <0.05<0.05 are marked with an asterisk. Note that these are uncorrected and none would remain significant under Bonferroni correction.
Dataset Augmentation TCN mAtt LaBraM
TUAB AmplitudeScaling 0.047 (p=0.746) 0.079 (p=0.590) -0.141 (p=0.335)
FrequencyShift 0.053 (p=0.717) 0.174 (p=0.232) 0.049 (p=0.738)
PhaseRandomisation 0.083 (p=0.571) -0.076 (p=0.609) 0.045 (p=0.761)
CAUEEG AmplitudeScaling -0.311 (p=0.095) -0.195 (p=0.302) 0.128 (p=0.499)
FrequencyShift -0.430* (p=0.018) 0.055 (p=0.772) 0.027 (p=0.885)
PhaseRandomisation -0.407* (p=0.026) -0.324 (p=0.081) -0.215 (p=0.253)
PhysioNet AmplitudeScaling 0.245 (p=0.155) -0.339* (p=0.046) -0.024 (p=0.886)
FrequencyShift -0.076 (p=0.663) -0.037 (p=0.833) 0.289 (p=0.075)
PhaseRandomisation -0.322 (p=0.059) -0.144 (p=0.409) 0.117 (p=0.503)
A.10 Pre-training Results on CAUEEG and PhysioNet
@@
A.11 Scaling Behaviour Plots with Pre-trained LaBraM Model
Figure 12: Scaling behaviour with pre-trained LaBraM model. Lines show the average accuracy across seeds with the standard error of the mean. Pre-training consistently improved LaBraM’s performance and allowed it to achieve comparable or even superior results to the other models. The lower performance on the PhysioNet dataset is due to differences in preprocessing (see Section˜A.6).